Find by use of differentials the approximate total area of a right circular cone whose base radius and height are 5.03 inches and 11.89 inches, respectively.

1
Expert's answer
2022-01-31T16:24:01-0500

A= "\pi"r(r+√h2+r2)

A= "\pi"rl+πr2

l=r2+h2

A= π·5.03·(5.03+11.892+5.032)

A= 283.49467